1) What volume of 2.25M H2SO4 should be diluted with water to form a 1.25L of 1.50M acid?


2) What is the concentration of the original solution if 450 mL; a certain solution is diluted in 750 mL water to form 2.0L of 1.50M HCl?


1
Expert's answer
2022-03-21T18:00:04-0400

1)If 1.5M = 1.25L

2.25M= ?

(2.25/1.5)× 1.25= 1.875L

2)n(HCl) = C(HCl) * V(solution) = 1.50 * 2.0 = 3.0 moles;

C0(HCl) = n(HCl)/V0 (solution) = 3/0.45 = 6.67 М.